Marco throws a rock straight down from a point 1760 ft above ground, with an initial velocity of 17 ft/sec.
A. Find an equation that models the height of the rock T seconds after it is launched, and graph the equation.
B. What is the maximum height of the rock?
C. How many seconds until it reaches the ground?

Answer by nerdybill(7384)

Marco throws a rock straight down from a point 1760 ft above ground, with an initial velocity of 17 ft/sec.
A. Find an equation that models the height of the rock T seconds after it is launched, and graph the equation.
h(t) = -32t^2 + 17x + 1760
.
B. What is the maximum height of the rock?
max height is at the vertex
t = -b/(2a)
t = -17/(2*(-32)) = 0.265625 secs
h(0.265625) = -32(0.265625)^2 + 17(0.265625) + 1760
h(0.265625) = 1762.26 feet
.
C. How many seconds until it reaches the ground?
set h(t) to zero solve for t:
h(t) = -32t^2 + 17x + 1760
0 = -32t^2 + 17x + 1760
applying quadratic equation yields:
t = {-7.16, 7.69}
throw out the negative solution leaving
t = 7.69 secs